# Rotate, Clip, and Partition: Towards W2A4KV4 Quantization by Integrating Rotation and Learnable Non-uniform Quantizer

### v1

@@ A.5 GEMM Kernel Design for Non-uniform W2A4 Quantization
A.6 Details and More Results on GEMV
@@
Ours 76.55 53.55 65.90 39.33 58.83
Ours 76.71 53.88 65.43 41.04 59.27
@@
Ours 75.35 53.95 67.64 41.80 59.69
Ours 76.16 54.35 71.19 42.75 61.11
Table 8: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks on LLaMA models with parameter sizes of at least 7B.
Model #Bits (W-A-KV) Configuration PIQA Hella. Wino. ARC-c Avg.
Method Rotation LDP
1B 16-16-16 75.30 60.70 60.90 38.70 58.90
2-4-16 BitDistiller 51.95 27.41 48.46 19.45 36.82
Ours 61.15 30.66 50.67 21.84 41.08
Ours 61.42 31.55 51.78 20.65 41.08
2-4-4 BitDistiller 55.33 26.62 48.46 19.79 37.55
Ours 61.75 30.05 51.22 20.05 40.77
Ours 60.71 31.54 53.51 21.42 41.80
3-4-16 BitDistiller 53.53 28.35 48.61 19.62 37.53
Ours 69.53 40.31 55.40 26.27 47.88
Ours 69.64 40.57 56.12 26.37 48.18
3-4-4 BitDistiller 54.18 28.26 50.90 21.67 38.75
Ours 68.98 37.80 55.40 26.36 47.14
Ours 68.12 39.30 56.12 26.11 47.41
3B 16-16-16 76.00 71.00 66.60 47.60 65.30
2-4-16 BitDistiller 54.02 26.80 52.48 18.25 37.89
Ours 65.99 36.51 52.48 26.19 45.29
Ours 65.43 37.35 54.70 25.43 45.71
2-4-4 BitDistiller 51.84 26.70 51.38 19.11 37.26
Ours 64.30 36.26 51.38 25.08 44.26
Ours 65.45 36.66 53.75 26.37 45.56
3-4-16 BitDistiller 52.72 26.66 50.43 19.45 37.32
Ours 74.04 49.56 63.22 35.83 55.66
Ours 73.77 49.52 62.65 37.54 55.87
3-4-4 BitDistiller 53.91 26.82 48.03 20.30 37.27
Ours 74.31 49.19 60.06 36.77 55.08
Ours 73.18 48.87 62.43 36.01 55.12
Table 9: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks on LLaMA-3.2 models with parameter sizes of at most 3B.
@@
Figure 8: A reasoning task example from HumanEval  Chen et al. (2021) benchmark, conducted by two differently quantized WizardCoder-7B  Luo et al. (2023) models. The results in the orange box is from state-of-the-art QAT method BitDistiller Du et al. (2024) with applying rotation. In the green box, our proposed RCP is applied. Both methods employ exactly the same 4-bit quantization setting for activation and KV-cache.
A.7 Information About Use of AI Assistants

### v2

@@ A.6 GEMM Kernel Design for Non-uniform W2A4 Quantization
A.7 Details and More Results on GEMV
@@
Block Tiling
Each thread block consists of 128 threads (4 warps), and we only tile along the output dimension and define the tile size as BH. The reason we do not follow the traditional 2-dimensional tiling is that both the input tokens and weights are stored in row-major format and have sub-byte packing along the column direction, which makes it hard to efficiently use high-bandwidth memory that performs best when reading 128B data consecutively. Also, global loads with small transactions and repeated shared stores complicate the pipeline design for latency hiding and degrade overall performance.
@@
Figure 8: Additional two-dimensional histogram plots with $p_{platy}=P(Kurt(\mathbf{W})<0)$ specified in each title.
@@
BitDistiller 76.55 53.55 65.90 39.33 58.83
RCP 76.71 53.88 65.43 41.04 59.27
2-13B 16-16-16 79.16 60.13 72.14 48.12 64.89
2-4-16 BitDistiller 61.86 33.40 53.51 23.46 43.06
BitDistiller 72.14 44.77 59.67 35.84 53.11
RCP 73.55 49.94 63.14 34.64 55.32
2-4-4 BitDistiller 57.45 30.73 50.35 20.39 39.73
BitDistiller 67.68 41.58 54.62 29.69 48.39
RCP 71.65 43.79 57.30 32.68 51.36
3-4-16 BitDistiller 75.29 53.91 62.50 38.56 57.57
BitDistiller 77.09 56.53 70.24 44.19 62.01
RCP 77.69 57.67 70.86 45.56 62.95
3-4-4 BitDistiller 75.68 49.94 64.00 39.50 58.07
BitDistiller 76.71 57.11 68.03 44.19 61.51
RCP 77.42 56.13 69.46 42.66 61.42
Table 13: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks on LLaMA models.
Model #Bits (W-A-KV) Configuration PIQA Hella. Wino. ARC-c Avg.
Method Rotation LDP
3.2-1B 16-16-16 75.30 60.70 60.90 38.70 58.90
2-4-16 BitDistiller 51.95 27.41 48.46 19.45 36.82
BitDistiller 61.15 30.66 50.67 21.84 41.08
RCP 61.42 31.55 51.78 20.65 41.08
2-4-4 BitDistiller 55.33 26.62 48.46 19.79 37.55
BitDistiller 61.75 30.05 51.22 20.05 40.77
RCP 60.71 31.54 53.51 21.42 41.80
3-4-16 BitDistiller 53.53 28.35 48.61 19.62 37.53
BitDistiller 69.53 40.31 55.40 26.27 47.88
RCP 69.64 40.57 56.12 26.37 48.18
3-4-4 BitDistiller 54.18 28.26 50.90 21.67 38.75
BitDistiller 68.98 37.80 55.40 26.36 47.14
RCP 68.12 39.30 56.12 26.11 47.41
3.2-3B 16-16-16 76.00 71.00 66.60 47.60 65.30
2-4-16 BitDistiller 54.02 26.80 52.48 18.25 37.89
BitDistiller 65.99 36.51 52.48 26.19 45.29
RCP 65.43 37.35 54.70 25.43 45.71
2-4-4 BitDistiller 51.84 26.70 51.38 19.11 37.26
BitDistiller 64.30 36.26 51.38 25.08 44.26
RCP 65.45 36.66 53.75 26.37 45.56
3-4-16 BitDistiller 52.72 26.66 50.43 19.45 37.32
BitDistiller 74.04 49.56 63.22 35.83 55.66
RCP 73.77 49.52 62.65 37.54 55.87
3-4-4 BitDistiller 53.91 26.82 48.03 20.30 37.27
BitDistiller 74.31 49.19 60.06 36.77 55.08
RCP 73.18 48.87 62.43 36.01 55.12
@@
BitDistiller 75.35 53.95 67.64 41.80 59.69
RCP 76.16 54.35 71.19 42.75 61.11
Table 14: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks on LLaMA models.
hotpotqa mqa_en triviaqa 2wikimqa musique samsum passage_count Avg.
FP16 30.45 33.76 85.72 26.32 9.74 37.74 4.0 32.53
BitDistiller 2.95 11.09 7.28 4.42 2.03 1.97 0.86 4.37
RCP 10.42 26.73 41.77 17.48 4.12 33.31 1.43 19.32
Table 15: Performance comparison on the LongBench dataset. W2A4KV4 quantization is applied to the LLaMA-2-7B-chat-4k model.
hotpotqa mqa_en triviaqa 2wikimqa musique samsum passage_count Avg.
FP16 15.74 24.07 84.67 13.8 8.81 42.73 0.07 27.13
BitDisitller 2.18 9.72 11.82 5.18 1.09 5.65 0.48 5.16
RCP 5.97 13.58 33.07 10.81 2.1 19.34 1.13 12.29
Table 16: Performance comparison on the LongBench dataset. W2A4KV4 quantization is applied to the LLaMA-2-7B-Instruct-32k model.
Layer Size W16A4 W16A4+FP32Had W16A4+FP16Had RCP RCP+FP32Had RCP+FP16Had
(2048, 2048) 0.168 0.274 0.248 0.131 0.248 0.214
(2048, 8192) 0.327 0.387 0.348 0.143 0.240 0.218
(3072, 3072) 0.228 0.483 0.373 0.131 0.295 0.265
(3072, 8192) 0.526 0.773 0.661 0.140 0.294 0.271
(4096, 4096) 0.369 0.510 0.398 0.133 0.250 0.221
(4096, 11008) 0.866 1.014 0.902 0.143 0.250 0.223
(4096, 14336) 1.108 1.255 1.146 0.142 0.247 0.226
Table 17: GEMV latency for the value projection is measured with the overhead from activation quantization. The layer size is composed as (input channel, output channel). All latency numbers are in milliseconds.
@@
A.8 Reasoning Task Example: HumanEval
Figure 9: A reasoning task example from HumanEval  Chen et al. (2021) benchmark, conducted by two differently quantized WizardCoder 7B  Luo et al. (2023) models. The results in the orange box is from state-of-the-art QAT method BitDistiller Du et al. (2024) with applying rotation. In the green box, our proposed RCP is applied. Both methods employ exactly the same 4-bit quantization setting for activation and KV-cache.
Figure 10: Comparison of reasoning ouptut on GSM8K using MetaMath 7B under W2A4KV4.
A.9 Information About Use of AI Assistants